Percent Composition and Empirical Formula




1. The chemical analysis of aspirin indicates that the molecule is 60.00% carbon, 4.44% hydrogen,



and 35.56% oxygen. Determine the empirical formula for aspirin.




2. **Glucose is a simple sugar produced during photosynthesis. It has a molecular formula of C6H12O6.




a. Compute for the percent composition.




b. Determine the empirical formula.

1
Expert's answer
2022-03-13T15:20:42-0400

1 Empirical formula

C. H. O

Moles. 60÷12. 4.44÷1 35.56÷16

= 5. 4.44. 2.2225

Mole ratio=5÷2.2225:4.44÷2.225:2.225÷2.225


=2.25:2:1

=9:8:4

The empirical formula is C9H8O4


2(a) Percentage composition of C6H12O6

Carbon = (12×6÷180)×100=40%

Hydrogen= (12×1÷180)×100= 6.66%

Oxygen=(16×6÷180)×100 = 53.33%


(b) C:H:O = 6:12:6

Simplest ratio =1:2:1

Empirical formula is CH2O